A gas station sells regular gas for $2.05 per gallon and premium gas for $2.75 a gallon. At the end of a business day 280 gallons of gas were sold, and receipts totaled $637. How many gallons of each type of gas were sold?
Found 2 solutions by mananth, rfer:
Answer by mananth(16946)

x= Number of gallons regular
y= Number of gallons premium

1.00 x + 1.00 y = 280.00 .............1
Total value
2.05 x + 2.75 y = 637.00 .............2
Eliminate y
multiply (1)by -2.75
Multiply (2) by 1.00
-2.75 x -2.75 y = -770.00
2.05 x + 2.75 y = 637.00
Add the two equations
-0.70 x = -133.00
/ -0.70
x = 190.00
plug value of x in (1)
1.00 x + 1.00 y = 280.00
190.00 + y = 280.00
y = 280.00 -190.00
y = 90.00
x= 190.00 Number of gallons regular
y= 90.00 Number of gallons premium
m.ananth@hotmail.ca

Answer by rfer(16322)

r+p=280
r=280-p
-------------
2.05r+2.75p=637
2.05(280-p)+2.75p=637
574-2.05p+2.75p=637
0.70p=637-574
0.70p=63
p=63/0.70
p=90 gal premium
-------------
2.05r+2.75*90=637
2.05r=637-247.5
r=389.5/2.05
r=190 gal regular